pzll solve this
An invoice shows a net price of $178.72 after trade discounts of 30%, 10%, and 4% have been deducted.

a.
What was the list price of the goods? (Round your answer to 2 decimal places.)

Price $

b.
What single rate of trade discount would be equivalent to the discount series? (Round your answer to 2 decimal places.)

Single rate
%

Answer by addingup(3677)

100-(100*0.30)= 70
70-(70*0.10) = 63
63-(63*0.4) = 60.48% is the single rate of trade discount that would be equivalent to the discount series
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
178.72/0.6048 = 295.50 was the list price of the goods.
J